In the period of a year, an investor lost twice as much as he gained . If he originally had $10,000 and now has $8,000 , what amount did he gain and what did he lose?
Answer by stanbon(75887)

In the period of a year, an investor lost twice as much as he gained . If he originally had $10,000 and now has $8,000 , what amount did he gain and what did he lose?
-----------------------
Let his gain be "x".
Then his loss is "2x".
Equation:
10,000 + x -2x = 8000
-x = -2000
x = 2000 (his gain)
2x = 4000 ( his loss)
========================
Cheers,
Stan H.